# dynamics of specialization in neural modules under resource constraints

### v5

@@ Abstract
It has long been believed that the brain is highly modular both in terms of structure and function, although recent evidence has led some to question the extent of both types of modularity. We used artificial neural networks to test the hypothesis that structural modularity is sufficient to guarantee functional specialization, and find that in general, this doesn’t necessarily hold. We then systematically tested which features of the environment and network do lead to the emergence of specialization. We used a simple toy environment, task and network, allowing us precise control, and show that in this setup, several distinct measures of specialization give qualitatively similar results. We further find that in this setup (1) specialization can only emerge in environments where features of that environment are meaningfully separable, (2) specialization preferentially emerges when the network is strongly resource-constrained, and (3) these findings are qualitatively similar across the different variations of network architectures that we tested, but that the quantitative relationships depend on the precise architecture. Finally, we show that functional specialization varies dynamically across time, and demonstrate that these dynamics depend on both the timing and bandwidth of information flow in the network. We conclude that a static notion of specialization, based on structural modularity, is likely too simple a framework for understanding intelligence in situations of real-world complexity, from biology to brain-inspired neuromorphic systems. We propose that thoroughly stress testing candidate definitions of functional modularity in simplified scenarios before extending to more complex data, network models and electrophysiological recordings is likely to be a fruitful approach.
@@ 1 Introduction
To allow us to tease apart the factors influencing the relationship between structure and function, we designed flexible artificial neural networks performing tasks in a highly controlled synthetic environment. Generally, the link between structural and functional modularity is context-dependent and involves a complex and dynamic interplay of several internal and external variables. However, it is unclear if structural modularity is either necessary or sufficient for the emergence of specialization. We show here that it is not sufficient by exhibiting a case where even under strict structural modularity conditions, modules exhibit entangled functional behaviors. We then explore the space of architectures, resource constraints, and environmental structure in a more systematic way, and find sets of necessary constraints (within our constrained setup) for the emergence of specialized functions in the modules. Finally, by measuring specialization in networks unrolled in time, we find that specialization actually has complex temporal dynamics governed by the dynamics of the inputs and communication between modules. Refining our conceptual understanding of modularity and specialization in this abstract setting may lead to fresh insights into the role of modularity in the brain and in brain-inspired algorithms and neuromorphic hardware.
@@ 2.1 Environment: Data and Tasks
In this framework, recovering the decision variables from a stimulus is what constitutes the sub-tasks: the essential building blocks which can then be used to perform a more general global task. The precise form of stimulus can vary, representing different sets of irrelevant variables. Changing the nature and structure of this stimulus is what allows us to control how difficult it is to recover a decision variable, hence the difficulty of the sub-tasks. We use (1) pairs of written MNIST [Deng, 2012] digits (with each digit drawn from the same set), or (2) pairs of EMNIST letters [Cohen et al., 2017] (with each letter drawn from separated subsets). The global task is the precise computation to be performed after decoding the sub-tasks, and can also vary, allowing for control over the amount of cooperation needed by the agents. We train the networks on a parity-based choice:
@@ 2.2 Networks
For each input sample, the module with the highest overall "certainty" (largest value) takes the decision. More precisely, the global output layer of the network $\mathbf{r}^{out}$ is defined to be the output layer of the sub-network with the largest maximum value. In practice, that means that only the module having produced the largest overall value is being used in computing the resulting loss value, and that gradients are back-propagated to the other (non-deciding) module only via the (sparse) communication weights.
@@ 3.1 Moderate structural modularity is not a sufficient condition for specialization
We wish to investigate whether imposing structural modularity is enough to guarantee specialization. To test this idea in a simple and controllable setting, we use an architecture composed of two densely recurrent modules, each receiving its own input (image of an MNIST digit) and producing its own outputs (the separate pathway architecture from Fig. 1.B). The global task requires each module to recognize the digit presented to it, but also depends on the parity of the other digit. It can therefore be broken down into two sub-tasks (digit recognition) that are able to be solved by the modules individually, but where solving the global task requires the modules to share at least one bit of information from the solution of their sub-task (as it depends on whether or not the parity of the two digits is the same or different). Modules consist of 25 neurons, a regime where the network can solve the task, without being over-parameterized (a parameter choice we will understand in more detail in the next section). Then by varying the fraction of active connections $p$ in the communication layer between modules, we can directly control the structural modularity of the model. For dense intra-module connectivity and a proportion $p$ of active inter-module connections, the standard measure of structural modularity can be calculated as $Q={(1-p)}/{2(1+p)}$ (see Section 2.3). Test accuracy on this task increases with $p$ (unsurprisingly given that this means more parameters), but is always above chance and below saturation (Fig. 5).
@@
Figure 2: Levels of functional specialization of networks with (A) varying modularity $Q$, or equivalently (B) inter-module connectivity $p$, measured by three different metrics (columns). All metrics indicate a similar trend, sharply rising only at extreme levels of $Q$ modularity.
@@ 3.2 Environmental structure and resource constraints determine specialization
Environmental effects. In the previous section, we explicitly designed the data to have a perfectly separable structure. However, features are rarely perfectly independent in real environments (a horse is more likely to have a green background, while a car is more likely to have a gray background). To model these complex dependencies in a simple way, we vary the covariance matrix of the distributions of digits between sub-tasks, essentially modifying the probability of seeing the two digits being equal, $\mathcal{D}_{1}=\mathcal{D}_{2}$. We find that high covariance $c$ prevents specialization (Fig. 3 bottom left, or top right of each sub-figure), as knowing one hidden variable is already highly informative of the other. In Supplementary Fig. 6, we can also see that when the inputs to the two modules are drawn from separate datasets (as is the case when training with E-MNIST Section 2.1) modules show stronger specialization. This would be the case in a multimodal setting, such as audio-visual integration. It makes sense that when sub-tasks have different input distributions, this should promote specialization, since the knowledge of decoding one wouldn’t transfer as much to the other. Moreover, the computational capacity needed to decode both would be higher when sub-tasks are different in nature, changing some of the metabolic cost effects we discuss later on.
Metabolic cost constraints. To understand how resource constraints could shape the structure-function relationship of neural networks, we vary the number of elements that are costly to maintain or manufacture (in terms of both energy and space): neurons in a module ($n$) and synapses between modules (proportion $p$ of $n^{2}$ possible, or total number $p_{s}=pn^{2}$). We then measure the resulting specialization of the networks. Both constraints, when pushed towards their minimal values, lead to specialization arising, but the extent of which depends strongly on the architectural choices (as discussed in the next point). We can also see that the precise interactions between environmental ($c$) and metabolic ($n,p$) variables strongly depend on the choice of the model architecture (Fig. 3).
Architecture choices. The effect of architecture choice is more subtle. A few salient points include:
@@
    Input scheme: We find that having separate input pathways constitutes a strong structural prior, leading to higher functional modularity, whereas having a shared input pathway lowers it considerably (Fig. 3.A/B vs Fig. 3.C). This intuitively makes sense, as modules are forced to work with what they’ve got. When presented with different inputs, the only source of collapse of specialization comes from the information leakage introduced by the interconnections. As such, the main factor driving specialization, in that case, is the number of synapses $p_{s}$ (Fig. 3.A top left). Nonetheless, having a shared input does not prevent the appearance of a specialization pattern, albeit at a reduced scale. We find that in that case, the significant parameter is the size of modules, ie the number of neurons they’re composed of (Figure Fig. 3.C top left). Modules are forced to specialize when arriving at the threshold at which a single module does not have the computational capabilities to predict both digits (even when having both as input).
@@
    Output scheme: With separate outputs, the modules compete for which module determines the global readout through a max function. In that case we see a quantitatively lower specialization (Fig. 3.A vs Fig. 3.B). This could be explained by the decision-making dynamics of modules competing for the readout (not shown). In this structural configuration, networks often resort to having a main decision-making module. Although the task is designed to be solvable in a modular fashion, having such a module take all decisions means it’s able to predict both digits and thus displays less specialization, when it comes to our metrics at least. One could nevertheless argue that resorting to such a separation of function (decision-making vs context-providing modules) is in its own right a form of specialization, albeit one that is not picked up by our metrics. Finally, having a common readout seems to have a coordinating effect leading to modules being more specialized across the board, alleviating the necessity of extreme resource constraints to see the emergence of specialization.
@@ 3.3 Function specialization can vary dynamically
Figure 4: Specialization dynamics. A. Specialization of different parts of the network at different time-steps. The network unrolled through time is shown below, with inputs presented to modules at time $t=0$ at the left hand side, and subsequent time steps moving towards the right finishing in the readout layer at the final time step. In the plot above is shown the specialization of the corresponding module at a given time, where specialization close to 1 means full specialization on digit 0, close to -1 means full specialization on digit 1, and close to 0 means no specialization. The dark dashed lines in both the network diagram and the plot show the timing of inter-module communication. Each colored line represents networks with different inter-connection levels $p$. We vary the communication timings in the inputs, and find specialization collapses when communicating at high bandwidth ($p$ high). B. With dynamic noise in the inputs, specialization drops continually, especially in high noise setting. C. With stochasticity in the input dynamics (digits 0 and 1 turned on at $t_{0}$ and $t_{1}$ respectively, indicated by colored arrows), specialization dynamics closely follows the inputs’ dynamics. Networks showed with 1 (dark blue), 10 (teal) and 100 (yellow) active inter-module connections.
@@ 4 Discussion
Our second main result is that imposing some degree of structural modularity on a network is not, in general, sufficient to induce functional specialization, even in a toy setting seemingly ideal for the emergence of specialization. We conclude that (a) in machine learning and neuromorphic computing applications, if we wish to encourage the emergence of functional specialization, structural modularity is not sufficient, and (b) in neuroscience, if we observe structural modularity, we should not necessarily conclude that these modules will be functionally specialized. However, structural modularity may well be one component of a richer approach to understanding functional specialization, including more detailed architectures, different learning rules, training regimes and datasets. By systematically varying details of our toy network, we found that as a first step towards such a richer approach, introducing resource constraints was a reliable way to encourage stronger functional specialization. Intuitively this makes sense, as when computational resources are abundant there is no incentive to share or re-use these resources.
Having shown that the link from structure to function is not as straightforward as one might think, an exciting follow-up research question would be to investigate the link between functional and compositional aspects of modularity. This link has already proved harder to understand than expected: even when neural networks do attain functionally specialized organization they do not always make compositional use of these modules [Bahdanau et al., 2019, Csordás et al., 2021, Mittal et al., 2022].
@@
Our approach has been very handcrafted, with precise control over structural modularity and connectivity in the networks. However, other studies have taken a more emergent approach: [Achterberg et al., 2023, Liu et al., 2023] investigated whether modular properties can emerge directly from the first principle of minimizing connection costs. Spatially-embedded networks, regularized to minimize connections-costs while learning, do end up displaying modular and small-world features, but we note that both works had to introduce an additional regularization or optimization technique to see it emerge. Understanding if, and if so how, structural and functional modularity can emerge from purely low level and naturalistic principles outside of a controlled setup thus remains an open question. The search for such low level principles as a factor driving emergence of higher-levels concepts is nonetheless very promising. Another recent study shows how regularizing for energy efficiency could lead networks to organize into a predictive coding regime [Ali et al., 2022], and is a good example of how taking into account physical substrates and constraints can lead to a better understanding of networks’ function.
@@
We investigated relatively small networks, solving simple tasks. This enabled us to separate the multiple factors influencing modularity, and to thoroughly explore the parameter, architecture, and environment space. What’s more, the simplicity of our setup makes for an ideal proving ground for concepts of modularity: if the current static definition of specialization already shows its limits here, it would then seem too simple to describe complex and dynamic real-world systems. However, the opposite argument could also be made, suggesting that such a simple setup prevents us from seeing the emergence of functional specialization. Other works have taken more holistic approaches, studying both structural and functional modularity in trained networks at the same time [Filan et al., 2021, Hod et al., 2022, Casper et al., 2022]. A natural extension of this work could also take into account multiple modules and tasks, and introduce varying environments. Indeed, such environments have been shown to favor modular solutions [Kashtan et al., 2007, Ellefsen et al., 2015], and could also lead to different time-scales in specialization dynamics (life-long learning). Not only that, but it has been shown that multi-task learning can drive the emergence of modularity [Yang et al., 2020, Dobs et al., 2022], and as such would be interesting to investigate. In that context, the concept of systematic generalization is of crucial importance, and the ability for an agent to recombine and reuse previously acquired knowledge could prove decisive to its survival. Such an ability being better enabled through modular architectures [Goyal et al., 2020, Kirsch et al., 2018, Bahdanau et al., 2019], we can see how the necessity for generalization could also promote modular architectures to emerge (altough this has yet to be shown, as discussed above). Finally, we only looked at the limited learning rule of supervised learning with backpropagation, and other learning methods could potentially lead to different results, as suggested by [Bakhtiari et al., 2021].

### v6

@@ Abstract
The brain is structurally and functionally modular, although recent evidence has raised questions about the extent of both types of modularity. Using a simple, toy artificial neural network setup that allows for precise control, we find that structural modularity does not in general guarantee functional specialization (across multiple measures of specialization). Further, in this setup (1) specialization only emerges when features of the environment are meaningfully separable, (2) specialization preferentially emerges when the network is strongly resource-constrained, and (3) these findings are qualitatively similar across several different variations of network architectures. Finally, we show that functional specialization varies dynamically across time, and these dynamics depend on both the timing and bandwidth of information flow in the network. We conclude that a static notion of specialization is likely too simple a framework for understanding intelligence in situations of real-world complexity, from biology to brain-inspired neuromorphic systems.
@@ 1 Introduction
To allow us to tease apart the factors influencing the relationship between structure and function, we designed flexible artificial neural networks performing tasks in a highly controlled synthetic environment. Generally, the link between structural and functional modularity is context-dependent and involves a complex and dynamic interplay of several internal and external variables. However, it is unclear the extent to which structural modularity is important for the emergence of specialization through training. We show here a case where even under strict structural modularity conditions, modules exhibit entangled functional behaviors. We then explore the space of architectures, resource constraints, and environmental structure in a more systematic way, and find sets of necessary constraints (within our constrained setup) for the emergence of specialized functions in the modules. Finally, by measuring specialization in networks unrolled in time, we find that specialization actually has complex temporal dynamics governed by the dynamics of the inputs and communication between modules. Refining our conceptual understanding of modularity and specialization in this abstract setting may lead to fresh insights into the role of modularity in the brain and in brain-inspired algorithms and neuromorphic hardware.
@@ 2.1 Environment: Data and Tasks
In this framework, recovering the decision variables from a stimulus is what constitutes the sub-tasks: the essential building blocks which can then be used to perform a more general global-task. The precise form of stimulus can vary, representing different sets of irrelevant variables. Changing the nature and structure of this stimulus is what allows us to control how difficult it is to recover a decision variable, hence the difficulty of the sub-tasks. We use (1) pairs of written MNIST [Deng, 2012] digits (with each digit drawn from the same set), or (2) pairs of EMNIST letters [Cohen et al., 2017] (with each letter drawn from separated subsets). The global task is the precise computation to be performed after decoding the sub-tasks, and can also vary, allowing for control over the amount of cooperation needed by the agents. We train the networks on a parity-based choice:
@@ 2.2 Networks
For each input sample, the module with the highest overall "certainty" (largest value) takes the decision. More precisely, the global output layer of the network $\mathbf{r}^{out}$ is defined to be the output layer of the sub-network with the largest maximum value. In practice, that means that only the module having produced the largest overall value is being used in computing the resulting loss value, and that gradients are back-propagated to the other (non-deciding) module only via the (sparse) communication weights. This choice was made to minimize interactions between the modules other than the explicit sparse connectivity, ensuring the necessity of training the sparse inter-module weights.
@@ 3.1 Moderate structural modularity is not a sufficient condition for specialization
We wish to investigate whether imposing structural modularity is enough to guarantee specialization. To test this idea in a simple and controllable setting, we use an architecture composed of two densely recurrent modules, each receiving its own input (image of an MNIST digit) and producing its own outputs (the separate-pathway architecture from Fig. 1.B). The global task requires each module to recognize the digit presented to it, but also depends on the parity of the other digit. It can therefore be broken down into two sub-tasks (digit recognition) that are able to be solved by the modules individually, but where (idealising somewhat) solving the global task requires the modules to share just one bit of information from the solution of their sub-task (as it depends on whether or not the parity of the two digits is the same or different). Modules consist of 25 neurons, a regime where the network can solve the task, without being over-parameterized (a parameter choice we will understand in more detail in the next section). Then by varying the fraction of active connections $p$ in the communication layer between modules, we can directly control the structural modularity of the model. For dense intra-module connectivity and a proportion $p$ of active inter-module connections, the standard measure of structural modularity can be calculated as $Q={(1-p)}/{2(1+p)}$ (see Section 2.3). Test accuracy on this task increases with $p$ (unsurprisingly given that this means more parameters), but is always above chance and below saturation (Fig. 5).
@@
Figure 2: Levels of functional specialization of networks with (A) varying modularity $Q$, or equivalently (B) inter-module connectivity $p$, measured by three different metrics (columns). All metrics indicate a similar trend, sharply rising only at extreme levels of $Q$ modularity. Data are presented as mean values with a shaded standard error envelope.
@@ 3.2 Environmental structure and resource constraints determine specialization
First are environmental effects: In the previous section, we explicitly designed the data to have a perfectly separable structure. However, features are rarely perfectly independent in real environments (a horse is more likely to have a green background, while a car is more likely to have a gray background). To model these complex dependencies in a simple way, we vary the covariance matrix of the distributions of digits between sub-tasks, essentially modifying the probability of seeing the two digits being equal, $\mathcal{D}_{1}=\mathcal{D}_{2}$. We find that high covariance $c$ prevents specialization (Fig. 3 bottom left, or top right of each sub-figure), as knowing one hidden variable is already highly informative of the other. In Supplementary Fig. 7, we can also see that when the inputs to the two modules are drawn from separate datasets (as is the case when training with E-MNIST Section 2.1) modules show stronger specialization. This would be the case in a multimodal setting, such as audio-visual integration. It makes sense that when sub-tasks have different input distributions, this should promote specialization, since the knowledge of decoding one wouldn’t transfer as much to the other. Moreover, the computational capacity needed to decode both would be higher when sub-tasks are different in nature, changing some of the metabolic cost effects we discuss later on.
Second we look at metabolic cost constraints: To understand how resource constraints could shape the structure-function relationship of neural networks, we vary the number of elements that are costly to maintain or manufacture (in terms of both energy and space): neurons in a module ($n$) and synapses between modules (proportion $p$ of $n^{2}$ possible, or total number $p_{s}=pn^{2}$). We then measure the resulting specialization of the networks. Both constraints, when pushed towards their minimal values, lead to specialization arising, but the extent of which depends strongly on the architectural choices (as discussed in the next point). We can also see that the precise interactions between environmental ($c$) and metabolic ($n,p$) variables strongly depend on the choice of the model architecture (Fig. 3).
Third, we investigate architecture choices: The effect of architecture choice is more subtle. A few salient points include:
@@
    Input scheme: We find that having separate input pathways constitutes a strong structural prior, leading to higher functional modularity, whereas having a shared input pathway lowers it substantially (Fig. 3.A/B vs Fig. 3.C). This intuitively makes sense, as modules are forced to "work with what they’ve got". When presented with different inputs, the only source of collapse of specialization comes from the information leakage introduced by the interconnections. As such, the main factor driving specialization, in that case, is the number of synapses $p_{s}$ (Fig. 3.A top left). Nonetheless, having a shared input does not prevent the appearance of a specialization pattern, albeit at a reduced scale. We find that in that case, the significant parameter is the size of modules, ie the number of neurons they’re composed of (Fig. 3.C top left). Modules are forced to specialize when arriving at the threshold at which a single module does not have the computational capabilities to predict both digits (even when having both as input).
@@
    Output scheme: With separate outputs, the modules compete for which module determines the global readout through a max function. In that case we see a quantitatively lower specialization (Fig. 3.A vs Fig. 3.B). This could be explained by the decision-making dynamics of modules competing for the readout (a simplified demonstration of this effect is showed in Supplementary Fig. 6). In this structural configuration, networks often resort to having a main decision-making module. Although the task is designed to be solvable in a modular fashion, having such a module take all decisions means it’s able to predict both digits and thus displays less specialization, when it comes to our metrics at least. One could nevertheless argue that resorting to such a separation of function (decision-making vs context-providing modules) is in its own right a form of specialization, albeit one that is not picked up by our metrics. Finally, having a common readout seems to have a coordinating effect leading to modules being more specialized across the board, alleviating the necessity of extreme resource constraints to see the emergence of specialization.
@@ 3.3 Function specialization can vary dynamically
Figure 4: Specialization dynamics. A. Specialization of different parts of the network at different time-steps. The network unrolled through time is shown below, with inputs presented to modules at time $t=0$ at the left hand side, and subsequent time steps moving towards the right finishing in the readout layer at the final time step. In the plot above is shown the specialization of the corresponding module at a given time, where specialization close to 1 means full specialization on digit 0, close to -1 means full specialization on digit 1, and close to 0 means no specialization. The dark dashed lines in both the network diagram and the plot show the timing of inter-module communication. Each colored line represents networks with different inter-connection levels $p$. We vary the communication timings in the inputs, and find specialization collapses when communicating at high bandwidth ($p$ high). B. With dynamic noise in the inputs, specialization drops continually, especially in high noise setting. C. With stochasticity in the input dynamics (digits 0 and 1 turned on at $t_{0}$ and $t_{1}$ respectively, indicated by colored arrows), specialization dynamics closely follows the inputs’ dynamics. Networks showed with 1 (dark blue), 10 (teal) and 100 (yellow) active inter-module connections. Data are presented as mean values with a shaded standard error envelope.
@@ 4 Discussion
Our second main result is that imposing some degree of structural modularity on a network is not, in general, sufficient to induce functional specialization, even in a toy setting seemingly ideal for the emergence of specialization. We conclude that (a) in machine learning and neuromorphic computing applications, if we wish to encourage the emergence of functional specialization, structural modularity could prove insufficient, and (b) in neuroscience, if we observe structural modularity, we should not necessarily conclude that these modules will be functionally specialized. One limitation on this conclusion is that we have relied on the well-established $Q$-metric from network theory. This metric is widely used in connectomics research, but may not be the best measure of structural modularity (although we did not systematically investigate alternatives).
Structural modularity may well be one component of a richer approach to understanding functional specialization, including more complicated architectures, different learning rules, training regimes, datasets and neuron model (for example, a spiking neural network would have very different bandwidth to an equivalent artificial neural network). By systematically varying details of our toy network, we found that as a first step towards such a richer approach, introducing resource constraints was a reliable way to encourage stronger functional specialization. Intuitively this makes sense, as when computational resources are abundant there is no incentive to share or re-use these resources. Having shown that the link from structure to function is not as straightforward as one might think, an exciting follow-up research question would be to investigate the link between functional and compositional aspects of modularity. This link has already proved harder to understand than expected: even when neural networks do attain functionally specialized organization they do not always make compositional use of these modules [Bahdanau et al., 2019, Csordás et al., 2021, Mittal et al., 2022].
@@
Our approach has been very handcrafted, with precise control over structural modularity and connectivity in the networks. However, other studies have taken a more emergent approach: [Achterberg et al., 2023, Liu et al., 2023] investigated whether modular properties can emerge directly from the first principle of minimizing connection costs. Spatially-embedded networks, regularized to minimize connection costs while learning, do end up displaying modular and small-world features, but we note that both works had to introduce an additional regularization or optimization technique to see it emerge. Understanding if, and if so how, structural and functional modularity can emerge from purely low level and naturalistic principles outside of a controlled setup thus remains an open question. The search for such low level principles as a factor driving emergence of higher level concepts is nonetheless very promising. Another recent study shows how regularizing for energy efficiency could lead networks to organize into a predictive coding regime [Ali et al., 2022], and is a good example of how taking into account physical substrates and constraints can lead to a better understanding of networks’ function.
@@
We investigated relatively small networks, solving simple tasks. This enabled us to separate the multiple factors influencing modularity, and to thoroughly explore the parameter, architecture, and environment space. What’s more, the simplicity of our setup makes for an ideal proving ground for concepts of modularity: if the current static definition of specialization already shows its limits here, it would then seem too simple to describe complex and dynamic real-world systems. However, the opposite argument could also be made, suggesting that such a simple setup prevents us from seeing the emergence of functional specialization. Other works have taken more holistic approaches, studying both structural and functional modularity in trained networks at the same time [Filan et al., 2021, Hod et al., 2022, Casper et al., 2022]. Such an approach could reveal even further de-correlation of function and structure, as in [Hod et al., 2022, Casper et al., 2022] where function is shown to emerge even without modular structure. In that sense by focusing only our efforts, and the scope of our metrics, on the two structurally defined modules we could already be missing functionally specialized yet decentralized structures. Further extensions of this work could also take into account multiple modules and tasks, and introduce varying environments. Indeed, such environments have been shown to favor modular solutions [Kashtan et al., 2007, Ellefsen et al., 2015], and could also lead to different time-scales in specialization dynamics (life-long learning). Not only that, but it has been shown that multi-task learning can drive the emergence of modularity [Yang et al., 2020, Dobs et al., 2022], and as such would be interesting to investigate. In that context, the concept of systematic generalization is of crucial importance, and the ability for an agent to recombine and reuse previously acquired knowledge could prove decisive to its survival. Such an ability being better enabled through modular architectures [Goyal et al., 2020, Kirsch et al., 2018, Bahdanau et al., 2019], we can see how the necessity for generalization could also promote modular architectures to emerge (altough this has yet to be shown, as discussed above). Finally, we only looked at the limited learning rule of supervised learning with backpropagation, and other learning methods could potentially lead to different results, as suggested by [Bakhtiari et al., 2021].
@@
5 Data and Code Availability
Data availability : We provide all the data needed to reproduce the figures shown in this manuscript, and provide the code to reproduce said data in this repository. It is to be noted that this has been a long-running project with some experiment needing substantial computational resources (Section Section 3.2 especially). To this end, we would encourage interested readers to investigate the data available rather than trying to re-run the entire parameter search.
Code availability: The original code used to generate experiments and results from this work is available here. This being a long-running projet, the code repository is not in a particularly clean state, containing multiple abandoned ideas and unnecessary code. Because of this we recommend interested readers to consult this new repository, which is cleaner, faster and contains everything needed to reproduce all results shown here. The latter was built from scratch after we had a much better vision of the final paper, and can be use to experiment with modular networks and with our framework in general.
6 Aknowledgments
We thank the team of the Neural-Reckoning lab, especially Marcus Ghosh and colleagues at Imperial such as Maxence Faldor and Jeremy Pitt for fruitful insights, feedback and discussions.
7 Author Contributions
Both authors contributed equally to the work. G.B and D.G conceptualized the ideas, methodology and goals. G.B contributed most of the code and ran experiments. G.B and D.G wrote the manuscript.